OpenStudy (anonymous):

81 to the 1/4 is equal to 2 true or false?

OpenStudy (anonymous):

Let's see...81^(1/4) = 2 ?

OpenStudy (campbell_st):

well the question is saying the 4th root of 81 is 2 so to check evaluate 2^4 = 2 *2 *2 *2 now make the decision.

OpenStudy (anonymous):

False

OpenStudy (anonymous):

yes